OpenStudy (anonymous):

I have 10 Algebra questions that I must finish within the next hour or I do not get to attend my graduation tomorrow PLEASE HELP!!!!!

OpenStudy (anonymous):

LEFT SIDE of the attachment PLEASE HELP!

OpenStudy (anonymous):

[2X^{2}\] +7X+6=0 [2X^{2}\] +4X+3X+6=0 2X(X+2)+3(X+2)=0 2X(X+2)+3(X+2)=0 (2X+3)+(X+2)=0 X=-2 X=-1.5